# In Context Learning and Reasoning for Symbolic Regression with Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) are transformer-based machine learning models that have shown remarkable performance in tasks for which they were not explicitly trained. Here, we explore the potential of LLMs to perform symbolic regression — a machine-learning method for finding simple and accurate equations from datasets. We prompt GPT-4 to suggest expressions from data, which are then optimized and evaluated using external Python tools. These results are fed back to GPT-4, which proposes improved expressions while optimizing for complexity and loss. Using chain-of-thought prompting, we instruct GPT-4 to analyze the data, prior expressions, and the scientific context (expressed in natural language) for each problem before generating new expressions. We evaluated the workflow in rediscovery of five well-known scientific equations from experimental data, and on an additional dataset without a known equation. GPT-4 successfully rediscovered all five equations, and in general, performed better when prompted to use a scratchpad and consider scientific context. We also demonstrate how strategic prompting improves the model’s performance and how the natural language interface simplifies integrating theory with data. Although this approach does not outperform established SR programs where target equations are more complex, LLMs can nonetheless iterate toward improved solutions while following instructions and incorporating scientific context in natural language.
Keywords Symbolic Regression  $\cdot$ Large Lanugae Models  $\cdot$ GPT-4
@@ 1 Introduction
Symbolic regression methods mainly use genetic algorithms [1, 2, 3, 4] that generate random expressions from data, optimize their parameters, and evaluate their fitness with respect to the data through an iterative process until a fitness level or a specific number of iterations is reached. Other approaches include using Markov chain Monte Carlo (MCMC) sampling [5, 6], mixed integer nonlinear programming [7, 8, 9], greedy tree searches [10], pre-trained transformer-based models [11, 12], and sparse matrix algorithms [13, 14, 15]. These techniques are broadly geared towards accelerating equation search or efficient multi-objective optimization, but they do not integrate reasoning. Researchers have long explored ways to make these algorithms more informed by guiding the search space based on the context of the data [16, 17, 18, 19, 6, 20, 21, 22, 23, 24, 25, 26, 27, 9, 28], including using large language models integrated with genetic algorithms [29]. As scientific data is strongly tied to theory, encoding them in the program narrows the vast search space and can make the programs more effective.
@@
This work explores an approach to symbolic regression using large language models (LLMs) for equation discovery. LLMs are machine learning models adept at understanding and generating natural language. At its core, an LLM uses the transformer architecture—a neural network developed by Google that scales very effectively and allows the training of models on massive datasets [30]. The term “large" in the language model refers to the size and intricacy of the network, along with the dataset on which it was trained. Prior to GPT-3, natural language processing (NLP) tasks were solved by pretraining language models on vast text datasets and fine-tuning them for specific tasks. However, GPT-3 demonstrated that language models can excel at tasks using in-context learning without necessitating fine-tuning [31, 32]. LLMs are now commonly used for tasks like chat, code generation, summarization, translation, etc. -- and quite remarkably, these tasks can be effectively accomplished by using English language as model input without the need for machine learning expertise. We have firsthand experience with these models’ capabilities by interacting with AI chatbots on platforms like ChatGPT, Claude, or Gemini. LLMs have a wide understanding of the world from their training data and can even solve simple math problems expressed in natural language [33, 34, 35]. They are contributing significantly in education and research [36, 37, 38, 39] medicine [40, 41], physical and social sciences [42, 43, 44], as well as in legal [45, 46], business [47, 48], and entertainment [49, 50, 51, 52] sectors.
@@
Existing transformer-based symbolic regression programs [53, 11, 12] use models pretrained on large databases of synthetically-generated dataset/expression pairs, designed specifically for symbolic regression tasks. Thus, these approaches learn to pattern-match between datasets and math expressions, but they don’t employ iterations like those in genetic programming that optimize expressions for complexity and loss. In that respect, LLMs have been used to imitate evolutionary algorithms (EA) [54, 55, 56, 57, 58, 59, 60], and have specifically been applied to solve symbolic regression (SR) problems [56, 59]. Meyerson and coworkers [56] developed a workflow that performs genetic programming (mutation, crossover, etc.) through prompts in LLMs and tested symbolic regression in two ways: first, by using a language model in all evolutionary operators except the fitness measure, and second, by only using a language model in the initialization, crossover, and mutation operators. The second approach more closely resembles our work, but what we do is even simpler: we task GPT-4 to generate and/or transform expressions freely. Furthermore, while these approaches leverage “in-context” learning and don’t require pretraining [56, 59], the context employed is limited to a list of previously-obtained expressions. We propose to expand the context to include data, as well natural language descriptions of the scientific context of the problem. To effectively use this context, we anticipate that the LLM will perform better if given time for analysis [61, 62, 63]. Therefore, we also incorporate zero-shot chain-of-thought prompting with a scratchpad [64] to frame equation generation for symbolic regression as a reasoning problem [65] in the context of the data and free-form scientific information.
@@
Symbolic regression requires equation generation and precise fitting of numerical constants. Yang and coworkers [66] show that LLMs can perform linear regression, optimizing constants in math expressions via feedback loops, without showing the LLM the analytical form. They find that “LLM can often calculate the solution directly from the analytic form," however we find LLMs to be unreliable and inefficient for such tasks. Consequently, we interleave LLM-based optimization with gradient-based optimization, following a similar approach as [58], iteratively refining prompts for more accurate output. In our work, an LLM guides optimization of the symbolic structure of the math expressions, while SciPy performs numerical optimization of the constants.
This work aligns closely with two recent works [67, 68] in the literature in terms of methodology in that LLMs were used to generate expressions, and iterations were performed separately for optimization. Although there are some variations in the workflow and significant differences in the explored datasets, the unique contribution of this work lies in using data and context in natural language as model input, and implementing a scratchpad to record model output, which can help reveal if there was a case of “test set leakage.” Merler and coworkers [67] used multimodal LLMs with image and text inputs (without context and/or scratchpad), but their benchmark datasets does not reflect scientific data. Conversely, Shojaee and coworkers [68] specifically focused on scientific problems (without data and/or scratchpad), and while their expression generation utilized cleaner Python functions it is difficult to evaluate if GPT-4 has encountered these questions as they do not include data but present the task as a problem similar to college homework assignments.
@@ 2 Methods
In our initial tests, we asked GPT-4 to generate mathematical expressions from scientific data. We used a simple prompt to assess its capability and tested GPT-3.5-turbo and GPT-4 at varying temperatures (Figures 1). In all cases, the models produced expressions while “hallucinating” arbitrary coefficients. We revised the prompts to ask the LLMs to “show all steps" to gain insight into how the model selects these values [63]. In response, the output either provided Python code for optimization or, at times, mathematical steps for optimization.
@@
Therefore, we designed a workflow (Figure 2) where we task GPT-4 with suggesting expressions without fitting constants, and subsequently, we optimize the coefficients of the expressions using SciPy outside the LLM. A Python class takes in expressions, optimizes them, then calculates their complexity and mean squared error (MSE). These results are stored as a dictionary, the text of which is passed back to GPT-4 in a subsequent prompt asking to suggest better expressions. We initially evaluated GPT-3.5-turbo, but found it less reliable in following instructions than GPT-4, more frequently generating expressions that didn’t parse.
Figure 2: Workflow for using LLMs for SR. First, the dataset is sent to GPT-4 (with or without context) which is instructed to suggest expressions without optimizing parameters. The generated expressions are then parsed by Python and optimized using SciPy (Nelder-Mead [70] method with basin-hopping as the numerical optimizer [71]). Results for each expression are stored in a Python dictionary, and added to a list of results from prior expressions. The top performing expressions are sent as feedback to GPT-4, which is asked to suggest better expressions optimizing for both complexity and loss. The feedback loop is run for a set number of iterations.
We use two prompts: 1) Initial Prompt - where we input data and ask GPT-4 to suggest expressions, and 2) Iteration Prompt - where, along with data, we provide feedback in JSON format. We sort the expressions in descending order of mean squared error, then share this list as feedback, asking GPT-4 to suggest new equations optimizing for complexity and loss. In addition to this, we also include a system message to guide the behavior of the language model, setting the tone of the conversation. We do not use chat history or any advanced forms of memory [72, 73, 74, 75] in this workflow. The history of generated expressions is maintained externally and provided as feedback in the iteration prompt. A Python function maintains this feedback loop by sorting and filtering a list of dictionaries based on MSE and complexity. At the start of the search, up to 6 expressions are always returned; later, the least accurate expressions that are not on the Pareto front are pruned in order to manage the length of the context window.
By making each call to GPT-4 independent, we aim to minimize hallucinations that have been observed when the chat history becomes too large in conversational models [76, 77], as well as manage cost by pruning the large quantity of generated scratchpad text.
@@
Removing bias: We aimed to make the workflow run smoothly without any human intervention and therefore, it was important to obtain machine-readable and precise output from GPT-4 to ensure the SciPy function runs without any errors. A simple way to illustrate the expected outcome was by providing examples in a few-shot context [31, 82]. While this led to expressions matching the required syntax, we noticed the generated expressions resembled the examples we provided (Figure 3). While this taught the LLM correct syntax, it introduced bias that severely compromised the search.
Figure 3: Bias from examples GPT-4 generated exact expressions from the prompt examples that were given to illustrate the output syntax. The revised prompt (Prompt 2) eliminates this bias from the equation prediction process.
This motivated our two-prompt setup, with an initial prompt tasking GPT-4 to generate unbiased expressions in LaTeX (which it generated more reliably than as Python strings), and an iteration prompt receiving previously-generated examples as Python script (now without bias). To bridge these, we converted the LaTeX text into SciPy-compatible text using a Python function for string formatting, which we developed after identifying the types of expressions GPT-4 was prone to generating. Although this approach did not completely resolve syntax errors, it effectively managed them for the GPT-4 model.
Recording analysis in a scratchpad: Studies have shown that LLM performance can be improved by slowing down the model or breaking down its tasks into smaller steps [61, 62]. One popular strategy is the “scratch pad" technique [64], which mimics how we solve problems by jotting down notes before presenting a final answer in exams. We implemented this in our workflow, instructing GPT-4 to generate responses in two parts: data analysis and observations in a scratch pad, followed by its conclusions. After implementing this technique, the model immediately generated higher-quality expressions (Figure 4).
@@
Preventing redundant expressions: GPT-4 often generated expressions like $x+c_{1}$ and $x-c_{1}$, implying they are different. However, since the constants are yet to be fitted, these expressions are the same from a symbolic regression perspective. While a computer algebra system like SymPy [83] could in principle catch some redundant expressions by simplification to a canonical form [6, 28], this wouldn’t distinguish “SR-similar" expressions that become equivalent after fitting constants. Instead, we used prompt engineering to guide generation toward unique expressions: we added a note in the iteration prompt with examples showing how expressions in symbolic regression are similar before parameters are optimized. While this didn’t completely resolve the issue, we did observe a reduction in occurrences, and at times, the scratchpad revealed GPT-4 correctly addressing this by taking these examples into account (see Figure 5).
@@
Avoiding uninteresting expressions: During the iterative runs, GPT-4 attempted to improve its accuracy by repeatedly adding linear terms to suggested expressions from previous iterations. To address this issue, we encouraged the model to explore diverse expressions in the prompt. Additionally, in cases involving datasets with multiple independent variables, GPT-4 sometimes recommended excluding variables that exhibited weak correlation with the overall dataset pattern. While this may be useful in some contexts, we wanted expressions that made use of all of the available data, so we explicitly instructed the use of all variables. Additional constraints we implemented included limiting the types of math operators to include and preventing generation of implicit functions, as shown in Figure 6.
@@
Classical SR approaches always incorporate some amount of guidance from the practitioners (e.g. by limiting available math operators and variables and incorporating some bias toward parsimonious expressions); more recent work enables SR programs to explicitly account for limiting behavior (shape-constrained SR) and dimensional constraints [84, 85, 86, 24, 18, 25, 87]. Incorporating these into SR algorithms typically requires bespoke modification of research software; our strategy is to use LLM prompting in natural language to instill this context, which might additionally include more fluid constraints like “generate diverse equations" and “consider scientific context." For instance, the context we provide for Kepler’s Law is a single line text stating, “The data is about planetary motion in astrophysics where the independent variable (x1) is semi-major axis, and the dependent variable (y) is period in days." Figure 11 in SI lists the context provided for all the datasets.
@@ 3 Results
We evaluate our workflow using experimental datasets associated with meaningful scientific context. SR benchmarks often use synthetic data; we designed our tests around a benchmark [88] specifically curated for evaluating SR algorithms for scientific data. From here, we selected three experimental datasets from astronomical observations (Bode’s Law, Hubble’s Law, and Kepler’s Law). To these, we added two experimental chemistry datasets: Langmuir and Dual-site Langmuir adsorption isotherm models [9].
Figure 12 in SI illustrates all the datasets with the target expression for each. We also tested our workflow using a dataset on friction losses in pipe flow [89]. This phenomena doesn’t have an established target model, and has been the subject of prior study by other SR programs [6]. We included the whole dataset for each problem in the prompt to GPT-4 at the start and in each iteration. Because the Hubble and Leavitt’s datasets had an unreasonable number of digits following each entry, we rounded to 3 decimal places when sending data to GPT-4 but used the original dataset when optimizing constants in SciPy. This reduced the tokens (and cost) of running GPT-4. We pass the entire dataset SR is different than other machine learning models, as we feed the entire dataset for training where the test is the output expression [88].
We conducted eight sets of experiments on each of the six datasets, running each 5 times to evaluate robustness. In four of the experiments, we used basic binary operators ($+$, --, *, $\div$), incorporating ‘sqrt’ for Kepler’s data and ‘^’ and ‘exp’ for Bode’s data. Empirical relations often involve field-specific operators, so the equation search should account for this. We refer to this set of tests as an “easy search" because the space of possible equations is constrained to that generated by the operators in our dataset. In addition, we conducted four further tests adding common unary operators (sqrt, log, exp, square, cube) alongside the basic ones to evaluate a more difficult search.
@@
Figure 7: GPT-4 results on “easy" and “hard" searches. Easy search includes only basic operators while harder one adds more. All the datasets were run in four settings: without context, without data, without scratchpad and finally with all tools on. The score at each iteration is a total of x/5 independent runs where target model expressions were found. With these settings; dual-site Langmuir was never found after 50 iterations; with a modified feedback loop, it was only found 1/5 times in easy searches with and without tools.
@@
Hubble: Hubble’s Law is represented by the simple equation y = c1*x, which is often a first guess when few operators are made available. However, including context plays a role; this is especially apparent in the “hard” search, where removing the scratchpad for reasoning or the context inhibited performance. This dataset is particularly noisy, as well; including the noisy data as context actually inhibits the search.
@@
Bode: The results for Bode’s Law indicate that, similar to other SR programs [88], GPT-4 encounters difficulty accurately rediscovering the target model expression. Curiously, it performs best when the context is excluded and worst when the data is excluded. This suggests that the context we provided was counterproductive for the search, and purely reasoning about the data led greater success. We also note that GPT-4 finds the target model as — $c_{1}\times\text{exp}(c_{2}\times x_{1})+c_{3}$ when all operators are provided, which is a symbolically-equivalent way of expressing Bode’s Law, though lacking the interpretability of the original form.
@@
Kepler: We find strong evidence that the prompt has triggered GPT-4’s memorization of Kepler’s Law: the scratchpad reveals GPT-4 associates the variable names in the context with Kepler’s Law, and it not only guesses the right answer in the first iteration, it names Kepler’s Law in its justification. Perhaps because this relationship is routinely taught in high school and college physics courses, and thus likely to be more represented in GPT-4’s training data than the other relationships. The hard search for Kepler’s Law with all tools on has led GPT-4 to explore more complex expressions since the target model complexity is only 5 with MSE of 46.6886. GPT-4 also finds the target model in the form $c_{1}x_{1}^{c_{2}}$, with $c_{2}$ left as a constant for optimization instead of $\frac{3}{2}$. However, SciPy optimizes and fits $c_{2}$ with to a floating point power ($\approx 1.5$). We did not consider this to be a rediscovery of the target model as such expressions are not dimensionally consistent.
@@
Langmuir: Langmuir’s model is more obscure, and was almost never guessed in the first round. In easier searches, GPT-4 consistently found it within 15 iterations; inclusion of context and data seemed to improve performance. However, the hard search was much more challenging, with Langmuir only discovered once, in the case with all tools on.

### v2

@@ Abstract
Large Language Models (LLMs) are transformer-based machine learning models that have shown remarkable performance in tasks for which they were not explicitly trained. Here, we explore the potential of LLMs to perform symbolic regression — a machine-learning method for finding simple and accurate equations from datasets. We prompt GPT-4 and GPT-4o models to suggest expressions from data, which are then optimized and evaluated using external Python tools. These results are fed back to the models, which proposes improved expressions while optimizing for complexity and loss. Using chain-of-thought prompting, we instruct the models to analyze data, prior expressions, and the scientific context (expressed in natural language) for each problem before generating new expressions. We evaluated the workflow in rediscovery of five well-known scientific equations from experimental data, and on an additional dataset without a known equation. GPT-4 and GPT-4o successfully rediscovered all five equations and generally performed better when prompted to use a scratchpad and consider scientific context. GPT-4o model demonstrated improved reasoning with data patterns, particularly evident in the dual-site Langmuir and Nikuradse datasets. We demonstrate how strategic prompting improves the model’s performance and how the natural language interface simplifies integrating theory with data. We also observe how theory can sometimes offset noisy data and, in other cases, data can make up for poor context. Although this approach does not outperform established SR programs where target equations are more complex, LLMs can nonetheless iterate toward improved solutions while following instructions and incorporating scientific context in natural language.
Keywords scientific discovery  $\cdot$ GPT-4  $\cdot$ adsorption  $\cdot$ Nikuradse  $\cdot$ symbolic regression
@@ 1 Introduction
Symbolic regression methods mainly use genetic algorithms [1, 2, 3, 4] that generate random expressions from data, optimize their parameters, and evaluate their fitness with respect to the data through an iterative process until a fitness level or a specific number of iterations is reached. Other approaches include using Markov chain Monte Carlo (MCMC) sampling [5, 6], mixed integer nonlinear programming [7, 8, 9], greedy tree searches [10], pre-trained transformer-based models [11, 12], and sparse matrix algorithms [13, 14, 15]. These techniques are broadly geared towards accelerating equation search or efficient multi-objective optimization, but they do not integrate reasoning. Researchers have long explored ways to make these algorithms more informed by guiding the search space based on the context of the data [16, 17, 18, 19, 6, 20, 21, 22, 23, 24, 25, 26, 27, 9, 28], including using large language models integrated with genetic algorithms [29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. As scientific data is strongly tied to theory, encoding them in the program narrows the vast search space and can make the programs more effective.
@@
This work explores an approach to symbolic regression using large language models (LLMs) for equation discovery. LLMs are machine learning models adept at understanding and generating natural language. At its core, an LLM uses the transformer architecture—a neural network developed by Google that scales very effectively and allows the training of models on massive datasets [39]. The term “large” in the language model refers to the size and intricacy of the network, along with the dataset on which it was trained. Prior to GPT-3, natural language processing (NLP) tasks were solved by pretraining language models on vast text datasets and fine-tuning them for specific tasks. However, GPT-3 demonstrated that language models can excel at tasks using in-context learning without necessitating fine-tuning [40, 41]. LLMs are now commonly used for tasks like chat, code generation, summarization, translation, etc. -- and quite remarkably, these tasks can be effectively accomplished by using English language as model input without the need for machine learning expertise. We have firsthand experience with these models’ capabilities by interacting with AI chatbots on platforms like ChatGPT, Claude, or Gemini. LLMs have a wide understanding of the world from their training data and can even solve simple math problems expressed in natural language [42, 43, 44]. They are contributing significantly in education and research [45, 46, 47, 48] medicine [49, 50], physical and social sciences [51, 52, 53], as well as in legal [54, 55], business [56, 57], and entertainment [58, 59, 60, 61] sectors.
@@
Existing transformer-based symbolic regression programs [62, 11, 12] use models pretrained on large databases of synthetically-generated dataset/expression pairs, designed specifically for symbolic regression tasks. Thus, these approaches learn to pattern-match between datasets and math expressions, but they don’t employ iterations like those in genetic programming that optimize expressions for complexity and loss. In that respect, LLMs have been used to imitate evolutionary algorithms (EA) [29, 30, 31, 32, 33, 34, 35], and have specifically been applied to solve symbolic regression (SR) problems [31, 34]. Meyerson and coworkers [31] developed a workflow that performs genetic programming (mutation, crossover, etc.) through prompts in LLMs and tested symbolic regression in two ways: first, by using a language model in all evolutionary operators except the fitness measure, and second, by only using a language model in the initialization, crossover, and mutation operators. The second approach more closely resembles our work, but what we do is even simpler: we task LLMs to generate and/or transform expressions freely. Furthermore, while these approaches leverage “in-context” learning and don’t require pretraining [31, 34], the context employed is limited to a list of previously-obtained expressions. We propose to expand the context to include data, as well natural language descriptions of the scientific context of the problem. To effectively use this context, we anticipate that the LLM will perform better if given time for analysis [63, 64, 65]. Therefore, we also incorporate zero-shot chain-of-thought prompting with a scratchpad [66] to frame equation generation for symbolic regression as a reasoning problem [67] in the context of the data and free-form scientific information.
@@
Symbolic regression requires equation generation and precise fitting of numerical constants. Yang and coworkers [68] show that LLMs can perform linear regression, optimizing constants in math expressions via feedback loops, without showing the LLM the analytical form. They find that “LLM can often calculate the solution directly from the analytic form," however we find LLMs to be unreliable and inefficient for such tasks. Consequently, we interleave LLM-based optimization with gradient-based optimization, following a similar approach as [33], iteratively refining prompts for more accurate output. In our work, an LLM guides optimization of the symbolic structure of the math expressions, while SciPy performs numerical optimization of the constants. This approach aligns closely with two recent works [36, 37] in the literature where LLMs were used to generate expressions and optimizations were performed via iterations. Although there are some variations in the workflow and significant differences in the explored datasets, the unique contribution of this work lies in using both data and context in natural language as model input, and implementing a scratchpad to record raw model output as it solves the problem. This promotes informal reasoning about the data and context, and helps reveal “test set leakage.”
@@ 2 Methods
In our initial tests, we asked LLMs to generate mathematical expressions from scientific data. We used a simple prompt to assess its capability and tested GPT-3.5-turbo and GPT-4 at varying temperatures (Figures 1). In all cases, the models produced expressions while “hallucinating” arbitrary coefficients. We revised the prompts to ask the LLMs to “show all steps" to gain insight into how the model selects these values [65]. In response, the output either provided Python code for optimization or, at times, mathematical steps for optimization.
@@
Therefore, we designed a workflow (Figure 2) where we task GPT-4 and GPT-4o with suggesting expressions without fitting constants, and subsequently, we optimize the coefficients of the expressions using SciPy outside the LLM. A Python class takes in expressions, optimizes them, then calculates their complexity and mean squared error (MSE). These results are stored as a dictionary, the text of which is passed back to LLMs in a subsequent prompt asking to suggest better expressions. We initially evaluated GPT-3.5-turbo, but found it less reliable in following instructions than GPT-4 and GPT-4o, more frequently generating expressions that didn’t parse.
Figure 2: Workflow for using LLMs for SR. First, the dataset is sent to LLMs (with or without context) which is instructed to suggest expressions without optimizing parameters. The generated expressions are then parsed by Python and optimized using SciPy (Nelder-Mead [70] method with basin-hopping as the numerical optimizer [71]). Results for each expression are stored in a Python dictionary, and added to a list of results from prior expressions. The top performing expressions are sent as feedback to LLMs, which is asked to suggest better expressions optimizing for both complexity and loss. The feedback loop is run for a set number of iterations.
We use two prompts: 1) Initial Prompt - where we input data and ask GPT-4 and GPT-4o to suggest expressions, and 2) Iteration Prompt - where, along with data, we provide feedback in JSON format. We sort the expressions in descending order of mean squared error, then share this list as feedback, asking the LLMs to suggest new equations optimizing for complexity and loss. In addition to this, we also include a system message to guide the behavior of the language model, setting the tone of the conversation. We do not use chat history or any advanced forms of memory [72, 73, 74, 75] in this workflow. The history of generated expressions is maintained externally and provided as feedback in the iteration prompt. A Python function maintains this feedback loop by sorting and filtering a list of dictionaries based on MSE and complexity. At the start of the search, up to 6 expressions are always returned; later, the least accurate expressions that are not on the Pareto front are pruned in order to manage the length of the context window.
By making each call to LLMs independent, we aim to minimize hallucinations that have been observed when the chat history becomes too large in conversational models [76, 77], as well as manage cost by pruning the large quantity of generated scratchpad text.
@@
Removing bias: We aimed to make the workflow run smoothly without any human intervention and therefore, it was important to obtain machine-readable and precise output from LLMs to ensure the SciPy function runs without any errors. A simple way to illustrate the expected outcome was by providing examples in a few-shot context [40, 82]. While this led to expressions matching the required syntax, we noticed the generated expressions resembled the examples we provided (Figure 3). While this taught the LLM correct syntax, it introduced bias that severely compromised the search.
Figure 3: Bias from examples GPT-4 generated exact expressions from the prompt examples that were given to illustrate the output syntax. The revised prompts (Prompt 2 and 3) eliminate this bias from the equation prediction process.
This motivated our two-prompt setup, with an initial prompt tasking LLMs to generate unbiased expressions in LaTeX and an iteration prompt receiving previously-generated examples as Python script (now without bias). To bridge these, we converted the LaTeX text into SciPy-compatible text using a Python function for string formatting, which we developed after identifying the types of expressions GPT-4 was prone to generating. Although this approach did not completely resolve syntax errors, it effectively managed them for the GPT-4 model. While running the GPT-4o model, we realized that most syntax errors stemmed from incorrect LaTeX formatting. To address this, we made a minor adjustment by removing the term "LaTeX" and allowing the model to generate expressions as Python strings instead. This small change was was sufficient to remove bias and was also less prone to syntax errors. We nervetheless processed the generated Python expressions through our formatting function, and it worked seamlessly.
Recording analysis in a scratchpad: Studies have shown that LLM performance can be improved by slowing down the model or breaking down its tasks into smaller steps [63, 64]. One popular strategy is the “scratchpad" technique [66], which mimics how we solve problems by jotting down notes before presenting a final answer in exams. We implemented this in our workflow, instructing LLMs to generate responses in two parts: data analysis and observations in a scratchpad, followed by its conclusions. After implementing this technique, the model immediately generated higher-quality expressions (Figure 4).
@@
Preventing redundant expressions: The LLMs often generated expressions like $x+c_{1}$ and $x-c_{1}$, implying they are different. However, since the constants are yet to be fitted, these expressions are the same from a symbolic regression perspective. While a computer algebra system like SymPy [83] could in principle catch some redundant expressions by simplification to a canonical form [6, 28], this wouldn’t distinguish “SR-similar" expressions that become equivalent after fitting constants. Instead, we used prompt engineering to guide generation toward unique expressions: we added a note in the iteration prompt with examples showing how expressions in symbolic regression are similar before parameters are optimized. While this didn’t completely resolve the issue, we did observe a reduction in occurrences, and at times, the scratchpad revealed the models would correctly address this by taking the examples into account (see Figure 5).
@@
Avoiding uninteresting expressions: During the iterative runs, LLMs attempted to improve its accuracy by repeatedly adding linear terms to suggested expressions from previous iterations. To address this issue, we encouraged the model to explore diverse expressions in the prompt. Additionally, in cases involving datasets with multiple independent variables, LLMs would sometimes recommend excluding variables that exhibited weak correlation with the overall dataset pattern. While this may be useful in some contexts, we wanted expressions that made use of all of the available data, so we explicitly instructed the use of all variables. Additional constraints we implemented included limiting the types of math operators to include and preventing generation of implicit functions, as shown in Figure 6.
@@
Classical SR approaches always incorporate some amount of guidance from the practitioners (e.g. by limiting available math operators and variables and incorporating some bias toward parsimonious expressions); more recent work enables SR programs to explicitly account for limiting behavior (shape-constrained SR) and dimensional constraints [84, 85, 86, 24, 18, 25, 87]. Incorporating these into SR algorithms typically requires bespoke modification of research software; our strategy is to use LLM prompting in natural language to instill this context, which might additionally include more fluid constraints like “generate diverse equations" and “consider scientific context." For instance, the context we provide for Kepler’s Law is a single line text stating, “The data is about planetary motion in astrophysics where the independent variable (x1) is semi-major axis, and the dependent variable (y) is period in days." Figure 17 in SI lists the context provided for each of the datasets.
@@ 3 Results
We evaluated our workflow using experimental datasets associated with meaningful scientific context. SR benchmarks often use synthetic data; we designed our tests around a benchmark [88] specifically curated for evaluating SR algorithms for scientific data. From here, we selected three experimental datasets from astronomical observations (Bode’s Law, Hubble’s Law, and Kepler’s Law). To these, we added two experimental chemistry datasets: Langmuir and Dual-site Langmuir adsorption isotherm models [9].
Figure 18 in SI illustrates all the datasets with the target expression for each. We also tested our workflow using a dataset on friction losses in pipe flow [89]. This phenomena doesn’t have an established target model, and has been the subject of prior study by other SR programs [6]. We included the whole dataset for each problem in the prompt to LLMs at the start and in each iteration. Because the Hubble’s dataset has an unreasonable number of digits following each entry, we rounded to 3 decimal places when sending data to LLMs but used the original dataset when optimizing constants in SciPy. This reduced the tokens (and cost) of running the models. We pass the entire dataset SR is different than other machine learning models, as we feed the entire dataset for training where the test is the output expression [88].
We conducted eight sets of experiments for each model (GPT-4 and GPT-4o) on six datasets, running each set five times to evaluate its robustness. In four of the experiments, we used basic binary operators ($+$, --, *, $\div$), incorporating ‘sqrt’ for Kepler’s data and ‘^’ and ‘exp’ for Bode’s data. Empirical relations often involve field-specific operators, so the equation search should account for this. We refer to this set of tests as an “easy search" because the space of possible equations is constrained to that generated by the operators in our dataset. In addition, we conducted four further tests adding common unary operators (sqrt, log, exp, square, cube) alongside the basic ones to evaluate a more difficult search. All the datasets were run in four settings: without context, without data, without scratchpad and finally with all tools on. In the following text, we discuss each dataset individually and analyze their performance. The score at each iteration is a total of x/5 independent runs where target model expressions were found.
@@
Hubble: Hubble’s Law is represented by the simple equation y = c1*x, which was often a first guess in the GPT-4 model when few operators are made available. However, including context plays a role; this is especially apparent in the “hard” search, where removing the scratchpad for reasoning or context inhibited performance (Figure 7). This dataset is particularly noisy, and including the noisy data as context actually inhibits the search as observed by consistent scores of 5/5 from both models when they were run without the data. Furthermore, the scratchpad in GPT-4o reveals that when the model first guesses Hubble’s Law, it then tends to reason and deduce a more complex, non-linear model in context of the noisy data.
Figure 7: GPT-4 and GPT-4o results on Hubble’s data. In the GPT-4 model, all the trends except ‘Without Context’ overlap at score 5 for the “Easy Search” plot, and the trends ‘With All Tools’ and ‘Without Data’ overlap at score 4 for the “Hard Search”. In the GPT-4o model, the trends for ‘Without Data’ and ‘Without Context’ overlap at score 0 for the “Easy Search” plot, and all the trends except ‘Without Data’ overlap at score 0 for the “Hard Search”. Transparency and different marker styles are used to differentiate these trends.
@@
Bode: The results for Bode’s Law indicate that similar to other SR programs [88], LLMs encounter difficulty in accurately rediscovering the target model expression. Curiously, it performs best when the context is excluded and worst when the data is excluded. This suggests that our context was counterproductive for the search, and purely reasoning about the data led to greater success, with GPT-4o leading the performance (Figure 8). Scratchpad shows that GPT-4o refers to Bode in a few experiments, specifically in relation to the ’semi-major axis’ variable identity provided in the context, which may have been overly specific. However, it never suggested the target model expression in those instances. We also notice GPT-4o rediscovering the target model expression in the first few iterations from just the data patterns (Figure 9). However, both GPT-4 and GPT-4o find the target model as — $c_{1}\times\text{exp}(c_{2}\times x_{1})+c_{3}$, which is a symbolically-equivalent way of expressing Bode’s Law, though lacking the interpretability of the original form.
Figure 8: GPT-4 and GPT-4o results on Bode’s data. For the “Hard Search” in the GPT-4 model, all the trends except ‘Without Context’ overlap at a score of 0. Transparency and different marker styles are used to differentiate these trends.
Figure 9: GPT-4o reasoning on Bode’s data. The scratchpad snippet is from the first iteration in Run 4 with all tools on for hard search.
@@
Kepler: We find strong evidence that the prompt has triggered LLM’s memorization of Kepler’s Law: the scratchpad reveals that both GPT-4 and GPT-4o models associate the variable names and data in the context with Kepler’s Law, and it not only guesses the right answer in the first iteration, it names Kepler’s Law in its justification. Perhaps because this relationship is routinely taught in high school and college physics courses, and thus likely to be more represented in OpenAI’s training data than the other relationships. However, scratchpad also shows that although GPT-4o infers Kepler’s relationship from data and/or variable identity, it does not always suggest the target model expression but reasons with alternative expressions based on dataset patterns and feedback from prior suggestions. Both the models also often find the target model in the form $c_{1}x_{1}^{c_{2}}$, with $c_{2}$ left as a constant for optimization instead of $\frac{3}{2}$. However, SciPy optimizes and fits $c_{2}$ with a floating point power ($\approx 1.5$). We did not consider this a rediscovery of the target model as such expressions are not dimensionally consistent.
Figure 10: GPT-4 and GPT-4o results on Kepler’s Third Law data. For the “Hard Search” in the GPT-4o model, the trends for ‘Without Data’ and ‘With All Tools’ overlap at a score of 0. Transparency and different marker styles are used to differentiate these trends.
@@
Langmuir: Langmuir’s model is more obscure, and was almost never guessed in the first round. In easier searches, both GPT-4 and GPT-4o consistently found it within 15 iterations; inclusion of context and data seemed to improve performance. However, the hard search was much more challenging, with Langmuir discovered a few times, mainly with all tools on. An interesting observation about the GPT-4o model is that it references Langmuir and other adsorption models (Freundlich or BET), particularly for experiments with easy search and without data, where the model focuses more on the adsorption context. However, similar to other benchmark datasets, it does not always propose the target model expression; instead, it methodically analyzes the information it receives, including the context and feedback.
Figure 11: GPT-4 and GPT-4o results on Langmuir’s adsorption data. For the “Hard Search” in the GPT-4 model, all the trends except ‘With All Tools’ overlap at a score of 0. Transparency and different marker styles are used to differentiate these trends.